. A rigid tank contains 10 π‘˜π‘” of water at 90π‘œ 𝐢. If 8 π‘˜π‘” of the water is in the liquid form


and the rest is in vapour form, determine:


i. the pressure in the tank,


ii. the volume of the tank.

1
Expert's answer
2022-03-27T10:56:05-0400

"p=p_{sat}(T)=70.2~kPa,"

"v_f=0.001036~\frac{m^3}{kg},"

"v_g=2.3593~\frac{m^3}{kg},"

"V_f=m_fv_f=0.01~m^3,"

"V_g=m_gv_g=4.72~m^3,"

"V=V_f+V_g=4.73~m^3."